The eufrak package

Frank Mittelbach  Rainer Schöpf  Michael Downes

03 November 1996 Version 2.02

1 Introduction

This package was written originally by Frank Mittelbach and Rainer Schöpf; later it was moved into the AMS-L^{A}TEX distribution with only minor modifications.

It sets up some font shape definitions to use the Euler script symbols in math mode. These fonts are part of the AMS font package which can be found on many TEX servers. It is also directly available from the AMS and from TEX user groups.

The package does not require the amstex package.

\EuFrak To access the Euler Fraktur alphabet a ⟨math alphabet identifier⟩ called \mathfrak is provided. For example, the input

 \[ \mathfrak{A} \neq \mathcal{A} \]

will produce

𝔄 ≠ A

Here is a complete table of the beautiful letters drawn by Hermann Zapf:

𝔄 𝔅 ℭ 𝔇 𝔈 𝔉 𝔊 ℌ ℑ
𝔍 𝔎 𝔏 𝔐 𝔑 𝔒 𝔓 𝔔 ℜ
𝔖 𝔗 𝔘 𝔙 𝔚 𝔛 𝔜 ℨ

2 The Implementation

If the amsfonts package is already loaded, it doesn’t really make sense to load eufrak as well.

  \@ifpackageloaded{amsfonts}{%
    \PackageWarning{eufrak}{The eufrak package is redundant if the
      amsfonts package is used}%
    \def\EuFrak{\mathfrak}% for bulletproofing
    \endinput
  }{}%
  \NeedsTeXFormat{LaTeX2e}
  \ProvidesPackage{eufrak}
    [1996/11/03 v2.2b Euler Fraktur fonts]

The font shapes for the Euler Script medium and bold are defined in the amsfonts.fdd file which comes with the AMS font package. We repeat their default definition here for reference only.

 \DeclareFontFamily{U}{euf}{\skewchar\font'60}
 \DeclareFontShape{U}{euf}{m}{n}{
     <5> <6> <7> <8> <9> gen * eufm
    <10> <10.95> <12> <14.4> <17.28> <20.74> <24.88>eufm10}{}
 \DeclareFontShape{U}{euf}{b}{n}{
     <5> <6> <7> <8> <9> gen * eufb
    <10> <10.95> <12> <14.4> <17.28> <20.74> <24.88>eufb10}{}

\EuFrak Now we define the ⟨math alphabet identifier⟩ \EuFrak both for the normal and

the bold math version

  \DeclareMathAlphabet\EuFrak{U}{euf}{m}{n}
  \SetMathAlphabet\EuFrak{bold}{U}{euf}{b}{n}

The ‘psamsfonts’ option is intended to mean that the Y&Y/Blue Sky Research PostScript versions of the AMSFonts should be used. In that font set, the only .tfm files provided below size 10 are 5 and 7, rather than 5,6,7,8,9. See amsfonts.dtx for more discussion.

  \DeclareOption{psamsfonts}{\input{Ueuf57.fd}}

Here is a table describing the action of the eucal, euscript, and eufrak packages.

Package Option  Commands provided
eucal none \mathcal
eucal  [mathcal] \mathcal
eucal  [mathscr] \mathscr (\mathcal unchanged)
euscript none  \EuScript (obsolete)
euscript [mathcal] \mathcal

eufrak none \mathfrak (as well as obsolete \EuFrak for compatibility)

The preferred command name is \mathfrak, which for now just calls the old command name \EuFrak.

  \newcommand{\mathfrak}{\EuFrak}

Process the package options.

  \ProcessOptions

The usual \endinput to ensure that random garbage at the end of the file doesn’t get copied by docstrip.

  \endinput